(a) Solve the following initial value problem:

dy/dx = (y^2 − 4) / x^2

y(1) = 0

(b) Sketch the slope field in the square −4 <x< 4,−4 <y< 4, and draw several solution curves. Mark the solution curve corresponding to your solution.

(c) What is the long term behaviour of the solution from (a) as x → +∞?

Is it defined for all x?

(d) Find the only solution that satisfies lim(x→+∞) y(x) = 2, and explain why there are no other solutions with this property.
Answer #1

Page No Paga Pr er hut the vaue en. e + sleha prelal -421 4 YY 12 d21 umdapemad X 1 -u X - -3 -3 2. 2 -2 2 2 2 1 2019/9/23 4:Pageu Date Page No. 2 3 3 -S them d alunys ৫9 স = व म to pooitiso aualuamt Jalorgs +1 + -2 e LI + + log01979723 14:09Date: Page No + uE41092tL o elefineal auaul yeo uitn lim YCa1 4 g + hoat too Totheu etopauty u this Aalution 2019/9/23 14:10Page 12 C stntegsuation th aidle log +C +C +C +C + (+Kh Nous / ४-२. 2019/9/23 14:08Date Page 2 Page No. liuietung eqnO ly eauatian +1 (+ H HC1+c) e -+C) 4-1+c)=1og() EL+C)=og +C 2019/9/23 14:08 CI